OpenStudy (anonymous):

Given the sequence: {2, 8, 32, 128, 512…}, S12 = ???

OpenStudy (anonymous):

S12=2(4^12-1)/4-1 which simplifies to S12=2(1677215)/3. You must solve from here :)

OpenStudy (anonymous):

so it's 1118143

OpenStudy (anonymous):

yeppp

OpenStudy (anonymous):

OH wait!!! it's 11184810!!

OpenStudy (anonymous):

i did the (4^12) part wrong sorry

OpenStudy (anonymous):

what's it's then

OpenStudy (anonymous):

11,184,810